# VidEdit: Zero-shot and Spatially Aware Text-driven Video Editing

### v3

@@ 5 Conclusion & Discussion
We introduced VidEdit, a lightweight algorithm for zero-shot semantic video editing based on latent diffusion models. We have shown experimentally that this approach conserves more appearance information from the input video than other diffusion-based methods, leading to lighter edits. Nevertheless, the approach has a few limitations. Common with Kasten et al. (2021), the capacity of the MLP mapping networks decreases for complex videos involving rapid movements and very long-term videos. Since our method relies on the quality of such atlas representations, one possible way to expand the scope of possible video edits would be to strengthen and robustify the neural layered atlases construction approach. Finally, video editing raises several ethical challenges that we discuss here. Open-source diffusion models are trained on large amounts of data and inherit their biases. Therefore, the use of such models can raise ethical issues, whether the target text prompt is intentionally harmful or not. In addition, video editing algorithms could be used with harmful intent such as harassment or propagating fake news. To mitigate potential misuse, the Stable Diffusion model is released under a license focused on ethical and legal use, stating explicitly that users “must not distribute harmful, offensive, dehumanizing content or otherwise harmful representations of people or their environments, cultures, religions, etc. produced with the model weights”.
Appendix A Experimental Setup
A.1 Semantic Metrics
Prompt Consistency. $\mathcal{C}_{\text{\scriptsize{Prompt}}}$ measures the average CLIP similarity between a target text query and each video frame, for all edited videos. For a unique pair (image; caption) the CLIP similarity writes: $CLIPScore(I,C)=\max(100*cos(E_{I},E_{C}),0)$ with $E_{I}$ the visual CLIP embedding for an image $I$, and $E_{C}$ the textual CLIP embedding for a caption $C$. The prompt consistency score then writes :
$\mathcal{C}_{\text{\scriptsize{Prompt}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{i=1}^{F_{k}}CLIPScore(V_{i}^{k},C^{k})$
where $V_{i}^{k}$ is the $i$th frame of video $k$, $C^{k}$ the edited caption of video $k$, $F_{k}$ the number of frames in video $k$ and $N$ the number of edited videos.
Frame Accuracy. $\mathcal{A}_{\text{\scriptsize{Frame}}}$ corresponds to average percentage of edited frames that have a higher CLIP similarity with the target text query than with their source caption, for all edited videos.
Directional Similarity. $\mathcal{S}_{\text{\scriptsize{Dir}}}$ quantifies how closely the alterations made to an original image align with the changes between a source caption and a target caption. For a single edited frame the similarity score writes: $SIMScore(I_{t},I_{s},C_{t},C_{s})=100*cos(E_{I_{t}}-E_{I_{s}},E_{C_{t}}-E_{C_{s}})$ with $E_{I_{t}}$ and $E_{I_{s}}$ the visual CLIP embeddings of the edited image $I_{t}$ and source image $I_{s}$, and $E_{C_{t}}$ and $E_{C_{s}}$ the textual CLIP embeddings of the edited caption $C_{t}$ and source caption $C_{s}$. The global directional similarity score then writes:
$\mathcal{S}_{\text{\scriptsize{Dir}}}=\frac{1}{N}\sum_{k=1}^{N}\frac{1}{F_{k}}\sum_{i=1}^{F_{k}}SIMScore\left(\prescript{i}{}{V_{t}^{k}},\prescript{i}{}{V_{s}^{k}},C_{t}^{k},C_{s}^{k}\right)$
with $\prescript{i}{}{V_{t}^{k}}$ and $\prescript{i}{}{V_{s}^{k}}$ the $i$th frame of the $k$th edited video and its source counterpart respectively and $C_{t}^{k},C_{s}^{k}$ their associated captions.
Frame Consistency. $\mathcal{C}_{\text{\scriptsize{Frame}}}$ is used to assess the temporal coherence of a video. It measures the CLIP similarity between consecutive video frames.
A.2 Similarity Metrics
Regarding content preservation, we have chosen three metrics that operate on different feature spaces in order to capture a rich description of perceptual similarities. LPIPS Zhang et al. [2018] operates on the deep feature space of a VGG network and has been shown to match human perception well. HaarPSI Reisenhofer et al. [2018] operates on a wavelet space and shows the highest correlation with human opinion. Finally, PSNR assesses the distance between images directly at a pixel level. These metrics are extensively described in the literature and we refer to it for further details.
Appendix B Ablation Visualization
@@
Appendix C Additional Results
C.1 VidEdit samples
Source Prompt: "A couple of people riding a motorcycle down a road" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "trees" + "mountains" $\rightarrow$ "snowy trees" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "trees" $\rightarrow$ "a mountain lake" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "potted plant" $\rightarrow$ "a bouquet of roses" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "person" + "motorcycle" $\rightarrow$ "two golden statues riding a motorbike" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Source Prompt: "A man riding a kiteboard on top of a wave in the ocean" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "sea" + "mountains" + "sky" $\rightarrow$ "sea with mountains in the background in Van Gogh style" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "person" $\rightarrow$ "a santa" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "sea" + "mountain" + "sky" $\rightarrow$ "a fire",  "person" $\rightarrow$ "a fireman" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "sea" + "mountain" + "sky" $\rightarrow$ "the milky way",  "person" $\rightarrow$ "an astronaut" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ C.2 Baselines Comparison
Appendix D Blending Effect
@@
$\mathbb{A}_{\scriptsize{\text{Blended}}}=\mathbb{A}_{\scriptsize{\text{RGB}}}\odot\alpha+\mathbb{I}\odot(1-\alpha)$
with $\mathbb{A}_{\scriptsize{\text{RGB}}}$ the RGB channels of an atlas representation, $\mathbb{I}$ a fully white image and $\alpha$ the atlas’ opacity values.
@@
Appendix E Editing Time
Focusing on the interative part of editing [^3] Steps as DDIM inversion or LNA construction being considered as pre-processing steps. in which users are interested in [^4] Editing the atlas and reconstructing the video for atlas based methods. Simply inferring the model for other baselines., Figure 12 underlines the lightweight aspect of our method. The panel on the left shows that VidEdit can perform a large number of edits on a 70 frame long video in significantly less time than other appoaches. As an illustration, VidEdit demonstrates approximately 30 times faster editing capabilities compared to Text2Live, which is the second leading baseline in terms of editing capacities. On the other hand, as depicted in the right panel, the use of VidEdit becomes increasingly time-efficient compared to other baselines, as the number of frames to edit growths.

### v4

@@ 5 Conclusion & Discussion
We introduced VidEdit, a lightweight algorithm for zero-shot semantic video editing based on latent diffusion models. We have shown experimentally that this approach conserves more appearance information from the input video than other diffusion-based methods, leading to lighter edits. Nevertheless, the approach has a few limitations. Common with Kasten et al. (2021), the capacity of the MLP mapping networks decreases for complex videos involving rapid movements and very long-term videos. Since our method relies on the quality of such atlas representations, one possible way to expand the scope of possible video edits would be to strengthen and robustify the neural layered atlases construction approach.
Supplementary Material
Appendix A Ablation Visualization
@@
Appendix B Additional Results
B.1 VidEdit samples
Source Prompt: "A couple of people riding a motorcycle down a road" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "trees" + "mountains" $\rightarrow$ "snowy trees" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "trees" $\rightarrow$ "a mountain lake" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "potted plant" $\rightarrow$ "a bouquet of roses" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "person" + "motorcycle" $\rightarrow$ "two golden statues riding a motorbike" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Source Prompt: "A man riding a kiteboard on top of a wave in the ocean" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "sea" + "mountains" + "sky" $\rightarrow$ "sea with mountains, Van Gogh style" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "person" $\rightarrow$ "a santa" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "sea" + "mountain" + "sky" $\rightarrow$ "a fire",  "person" $\rightarrow$ "a fireman" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Target Edit: "sea" + "mountain" + "sky" $\rightarrow$ "the milky way",  "person" $\rightarrow$ "an astronaut" Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ B.2 Baselines Comparison
Appendix C Blending Effect
@@
$\mathbb{A}_{\text{Blended}}=\mathbb{A}_{\text{RGB}}\odot\alpha+\mathbb{I}\odot(1-\alpha)$
with $\mathbb{A}_{\text{RGB}}$ the RGB channels of an atlas representation, $\mathbb{I}$ a fully white image and $\alpha$ the atlas’ opacity values.
@@
Appendix D Atlas Construction
The atlas construction method takes as input a video and rough masks delineating the object(s) of interest. The objective is to compute (1) a collection of 2D atlases, one for the background and one for each dynamic object of interest; (2) a mapping from each pixel in the video to a 2D coordinate in each atlas; (3) opacity values at each pixel concerning each atlas. Each component is represented via coordinate-based MLPs. For the editing purpose, atlases are discretized into a fixed image grid $(1000\times 1000)$.
First, the mapping networks $\mathbb{M}_{b},\mathbb{M}_{f}$ receive a pixel location $p=(x,y,t)\in\mathbb{R}^{3}$ as input and output its corresponding 2D point $(u,v)\in\mathbb{R}^{2}$ in each atlas
$\mathbb{M}_{b}(p)=(u_{b}^{p},v_{b}^{p}),\qquad\mathbb{M}_{f}(p)=(u_{f}^{p},v_{f}^{p})$
The predicted 2D coordinates are then fed to an atlas network $\mathbb{A}$, that outputs the atlas’ RGB color at that location. While separate networks $\mathbb{A}_{f}$, $\mathbb{A}_{b}$ could be learned to represent foreground and background, it is sufficient to use a single atlas $\mathbb{A}$, and restrict mapping networks $\mathbb{M}_{b},\mathbb{M}_{f}$ to point into separate pre-defined quadrants in continuous [-1,1] space. The 2D atlas coordinates are then passed through a positional encoding denoted by $\phi(\cdot)$, to represent high frequency appearance information. The predicted colors are provided by:
$\mathbb{A}(\phi(u_{b}^{p}),\phi(v_{b}^{p}))=c_{b}^{p},\qquad\mathbb{A}(\phi(u_{f}^{p}),\phi(v_{f}^{p}))=c_{f}^{p}$
In addition, each pixel location is also fed into the alpha MLP, $\mathbb{M}_{\alpha}(\phi(p))=\alpha^{p}$ which outputs the opacity of each atlas at that location. The decomposition of the foreground and background layers is achieved by bootstrapping the alpha network using rough object masks that are computed with a pre-trained segmentor. Utilizing these networks, the reconstructed RGB color at each video pixel is estimated by alpha-blending the corresponding atlas colors such that
$c^{p}=(1-\alpha^{p})c_{b}^{p}+\alpha^{p}c_{f}^{p}$
This framework is trained end-to-end, in a self-supervised manner, where the main loss is a reconstruction loss to the original video. Additionally, regularization losses on the mapping and decomposition enforce the learning of a meaningful and semantic atlas that can be used for editing:
  1. 1.
    Rigidity loss: The local structure of objects is preserved as they appear in the input video by encouraging the mapping from video pixels to atlas to be locally rigid.
  2. 2.
    Consistency loss: Corresponding pixels in consecutive frames of the video are forced to be mapped to the same atlas point; pixel correspondence is computed using an off-the-shelf optical flow method.
  3. 3.
    Sparsity loss: Mapping networks are encouraged to recover the minimal content needed to recover the video in atlases via a sparsity loss.
The total loss is given by:
$\mathcal{L}=\mathcal{L}_{color}+\mathcal{L}_{rigid}+\mathcal{L}_{flow}+\mathcal{L}_{sparsity}$
Additional details about these loss terms can be found in (Kasten et al., 2021). We follow the implementation setup described in this paper to obtain the discretized atlases that are used to perform editing.